# QAEncoder: Towards Aligned Representation Learning in Question Answering System

### v1

@@ Abstract.
Modern QA systems entail retrieval-augmented generation (RAG) for accurate and trustworthy responses. However, the inherent gap between user queries and relevant documents hinders precise matching. Motivated by our conical distribution hypothesis, which posits that potential queries and documents form a cone-like structure in the embedding space, we introduce QAEncoder, a training-free approach to bridge this gap. Specifically, QAEncoder estimates the expectation of potential queries in the embedding space as a robust surrogate for the document embedding, and attaches document fingerprints to effectively distinguish these embeddings. Extensive experiments on fourteen embedding models across six languages and eight datasets validate QAEncoder’s alignment capability, which offers a plug-and-play solution that seamlessly integrates with existing RAG architectures and training-based methods. The code is available at https://github.com/IAAR-Shanghai/QAEncoder.
$\dagger$ The first two authors have equal contributions.
* Corresponding Author.
The first author completed this work during an internship at IAAR.
1. Introduction
“What I cannot create, I do not understand.”
— Richard Feynman
@@ 1. Introduction
Question Answering (QA) systems aim to generate accurate responses to user queries with applications in customer service (Xu et al., 2024), search engine (Ojokoh and Adebisi, 2018), healthcare (Guo et al., 2022) and education (Levonian et al., 2023), necessitating proficiencies in information retrieval, comprehension, and generation. Modern QA systems leverage large language models (LLMs) such as ChatGPT (Achiam et al., 2023), supplemented with retrieval-augmented generation (RAG) to address issues of outdated or hallucinatory information, especially for rapidly evolving knowledge bases (Lewis et al., 2020; Huang et al., 2023; Gupta et al., 2024). The efficacy of RAG hinges on its retrieval module for identifying relevant documents from a vast corpus. Dense retrievers (Lewis et al., 2020; Hofstätter et al., 2021), contrasted with keyword-matching-based sparse retrievers (Jones, 1973; Robertson and Zaragoza, 2009), have enabled efficient and precise retrieval by mapping queries and documents into a shared vector space. Despite advancements, a significant challenge that persists is bridging the semantic and syntactic gap between user queries and documents, known as the document-query gap (Zheng et al., 2020). Three main approaches have emerged to address this challenge: training-based alignment, document-centric alignment, and query-centric alignment.
Training-based approaches (Dong et al., 2022; Li et al., 2022; W et al., 2023; Zhang et al., 2024a; Khanna and Subedi, 2024) fine-tune or train embedding models from scratch with dedicated QA datasets to close the representation of relevant queries and documents, but struggle to generalize across new domains (Suprem and Pu, 2022). Furthermore, catastrophic forgetting necessitates updating embeddings for the previously encoded corpus when new data is learned (Pan et al., 2024), which is resource-intensive and practically unaffordable. Document-centric approaches (Wang et al., 2023a; Gao et al., 2023; Kim and Min, 2024) generate pseudo-documents for user queries by LLMs, which are then used to retrieve relevant information from the corpus, bridging the document-query gap by LLMs’ zero-shot generalization capabilities. However, pseudo-document generation is costy and time-consuming, and may include fabricated content (Wang et al., 2023a), increasing the risk of hallucinated information.
Query-centric methods  (Nogueira et al., 2019; Cheriton, 2019; Mallia et al., 2021a) generates and indexes document-relevant questions to align the indexed context with user queries in a training-free manner. However, existing solutions are predominantly based on sparse retrievers and have not fully leveraged the potential of dense retrievers. Combining query-centric methods with dense retrievers remains heavily under-explored. The naive solution can be storing all QA pairs into a vector database, but with several notable challenges:
  • C1.
    Expanded Index Size. Storing all QA pairs significantly increase the index size, leading to substantial expansion in storage requirements, especially problematic for large-scale corpora.
  • C2.
  • C3.
    Limited Query Handling. Although storing QA pairs individually can address predicted queries, this approach lacks robustness when confronted with the wide-ranging and diverse nature of potential queries. Such a method may be unable to effectively handle rephrasing in linguistic style, sentence structure, or even vocabulary, thereby compromising the overall reliability and performance of the system (Alting von Geusau and Bloem, 2020).
@@
Inspired by Feynman’s philosophy of learning (Reyes et al., 2021), we believe that effective information retrieval for QA systems extends beyond mere information storage and fundamentally depends on the active processing, involving the query formulation. This process mirrors human learning, where a deeper understanding of stored knowledge is achieved through thoughtful reflection and inquiry. For instance, the well-established 5W1H framework (Who, What, When, Where, Why, How) (Jinks, 2019) can be employed to systematically deconstruct information, fostering a comprehensive and nuanced understanding. In this paper, we continue the research line of query-centric methods and introduce an innovative approach called QAEncoder, motivated by a key observation, termed conical distribution hypothesis. Specifically, for a given document, its potential queries are embedded approximately within a single cluster on some hyperplane in the semantic space, while the document embedding lie on the perpendicular line intersecting the cluster center. Hence, the projection of the document embedding onto this hyperplane, i.e. the cluster center, is significantly closer to the potential queries, and can be utilized to optimize the original document embedding.
As demonstrated in Fig. 1, our method initially generates diversified queries (e.g. 5W1H) and then estimates the cluster center using Monte Carlo methods, i.e. mean pooling. The similarity matrix reveals that, compared to both document-query and query-query similarities, the mean-query similarity is significantly higher than both document-query and query-query similarities. Hence, we advocate using the cluster center as a surrogate for the document embedding in QA systems, which bridges the document-query gap robustly without extra index size and retrieval latency.
Despite these advantages, the basic implementation encounters a critical challenge. While enhancing similarity with user queries, it simultaneously reduces the distinguishability between document representations, as they all become query-like. To address this side effect, we further introduce document fingerprint strategies to infuse unique identities into document representations and enable state-of-the-art performance.
@@
The main contributions can be summarized as follows:
  • Methodological Innovations. We pioneer to bridge the document-query gap in dense retrievers from the query-centric perspective. Our method, QAEncoder, not only avoids extra index storage, retrieval latency, training cost and hallucination, but also guarantees diversified query handling and robust generalization. We further propose document fingerprint strategies to address the side effect of distinguishability and achieve state-of-the-art performance.
  • Theoretical Discovery. We formulate the conical distribution hypothesis, providing a theoretical foundation for the alignment of document and query embeddings, which is validated through extensive empirical analysis and provides deeper insights into semantic spaces in QA systems.
  • Practical Applications. QAEncoder is designed as a truly plug-and-play solution, seamlessly integrating with many existing retrieval and RAG architectures. This integration significantly enhances system performance with minimal modifications required.
Figure 1. Illustration of QAEncoder’s Semantic Alignment Process: Solid lines represent diversified query generation, while dashed lines indicate Monte Carlo estimation with embedding mean pooling. The heatmap depicts the similarity scores among the embeddings of the different queries, the document, and the mean estimation.
2. Related Works
@@
Retrieval-Augmented Generation (RAG) significantly improves large language models (LLMs) in question-answering (QA) systems by incorporating a retrieval module that fetches relevant information from external knowledge sources (Févry et al., 2020; Guu et al., 2020; Izacard and Grave, 2021; Zhao et al., 2024). Retrieval models have evolved from early sparse retrievers, such as TF-IDF (Jones, 1973) and BM25 (Robertson and Zaragoza, 2009), which rely on word statistics and inverted indices, to dense retrieval strategies (Lewis et al., 2020) that utilize neural representations for enhanced semantic matching . Advanced methods, such as Self-RAG (Asai et al., 2023) which determines if additional information is required and evaluates the relevance of retrieved content, and RAG-end2end (Siriwardhana et al., 2023) that jointly trains the retriever and generator, represent significant developments in this area. However, these methods still ignore the inherent semantic gap between queries and documents.
@@
Training-based approaches bridge the document-query gap generally by contrastive learning  (Xiong et al., [n.d.]; Qu et al., 2021) or knowledge distillation  (Zhang et al., 2024a; Khanna and Subedi, 2024). For instance,  (Dong et al., 2022) show parameter sharing of the query encoder and the document encoder improves overall performance by projecting queries and documents into shared space. Dual-Cross-Encoder (Li et al., 2022) and Query-as-context (W et al., 2023) train embedding models from scratch with paired document-query samples. E5 (Wang et al., 2024b) and GTE (Zhang et al., 2024b) concatenate different prompts before queries for instruction-tuned embedding models that enhance downstream task adaptability, especially for QA systems. However, training-based methods face high costs, generalization difficulties (Suprem and Pu, 2022), and catastrophic forgetting issue (Pan et al., 2024). Catastrophic forgetting, where the encoder continually trained on new data lose previously function, is particularly detrimental as it necessitates the costly re-encoding of entire corpora.
@@
Document-centric methods, such as HyDE (Gao et al., 2023) and Query2doc (Wang et al., 2023a), dynamically transform user queries into pseudo-documents using large language models for both sparse and dense retrievers. QA-RAG (Kim and Min, 2024) advances by implementing a two-way retrieval mechanism that utilizes both user query and pseudo-documents for respective retrieval. However, their effectiveness is highly dependent on the quality of pseudo-documents generated, which are susceptible to hallucinations, especially for latest information. Furthermore, invoking LLMs for each query imposes substantial computational costs and increases latency, leading to degraded user experience.
@@
The seminal work, Doc2Query (Nogueira et al., 2019), mainly focus on the vocabulary mismatch problem for sparse retrievers by expanding the document with keywords in predicted queries. The subsequent DocT5Query (Cheriton, 2019) improves by training a T5 model to predict queries. Based on DocT5Query, DeepImpact (Mallia et al., 2021b) further assigns weights on keywords with neural estimation for more precise sparse representations. Though these approaches have succeed for sparse retrievers, the integration with dense retrieval systems remains largely unexplored.
Figure 2. QAEncoder Architecture. Left: Corpus documents are embedded using QAEncoder to obtain query-aligned representations for indexing. User queries are encoded with a vanilla encoder and used to retrieve relevant documents. Right: Internal mechanism of QAEncoder. QAEncoder addresses the document-query gap by generating a diverse set of queries for each document to create semantically aligned embeddings. Additionally, document fingerprint strategies are employed to ensure document distinguishability.
3. Method
3.1. Problem Formulation
Given a query $q$ and a document corpus $\mathcal{D}=\{d_{1},d_{2},...,d_{i},...,d_{N}\}$, our task is to retrieve a subset of $K$ most relevant documents $\mathcal{D}_{+}=\{d{i_{1}},d_{i_{2}},...,d_{i_{j}},...,d_{i_{K}}\}$ through vector search. We define our embedding model as $\mathcal{E}(\cdot)$, which maps each document $d$ and query $q$ from the textual space $\mathcal{C}$ to a vector space $\mathbb{R}^{r}$. The semantic relevance is quantified by the cosine similarity between the query $q$ and each document $d$ in the embedding space, defined as:
(1) $sim(q,d)=\frac{\mathcal{E}(q)^{T}\mathcal{E}(d)}{\|\mathcal{E}(q)\|\|\mathcal{E}(d)\|}.$
Furthermore, for each document $d$ in our dataset, we invoke the query generator $\mathcal{Q}(\cdot)$ multiple times to generate $n$ predicted queries $\{q_{i}\}_{i=1}^{n}$, where the cluster center in embedding space is captured by $\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$ and estimated by Monte Carlo sampling $\overline{\mathcal{E}(\mathcal{Q}(d))}$.
@@ 3.1. Problem Formulation
Figure 3. Conical Distribution Hypothesis Validation. The figure presents three visualizations supporting the conical distribution hypothesis : (a): t-SNE visualization of queries derived from various documents in the embedding space, illustrating distinct clustering behavior. (b): Angular distribution of document and query embeddings, showing the distribution of angles between $v_{d}=\mathcal{E}(d)-\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$ and $v_{q_{i}}=\mathcal{E}(q_{i})-\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$. The angles form a bell curve just below 90°, supporting that $v_{d}$ is nearly orthogonal to each $v_{q_{i}}$ and serves as the normal vector to a hyperplane $\mathcal{H}$. (c): 3D visualization illustrating the conical distribution of document (black point) and query (colored points) embeddings within a unit sphere. The star indicates the queries’ cluster center.
3.2. Conical Distribution Hypothesis
@@
In this subsection, we formally define the highly simplified conical distribution hypothesis and validates its reasonableness with empirical analysis detailed in the appendix A.5.
@@
For any document $d$, the potential queries approximately form a single cluster on some hyperplane =$\mathcal{H}=\{x\in\mathbb{R}^{r}\mid w\cdot x=b\}$ in the semantic space, where $w\in\mathbb{R}^{r}$ is the normal vector and $b\in\mathbb{R}$ is the bias term. Furthermore, the document embedding $\mathcal{E}(d)$ lies on the perpendicular line intersecting the cluster center $\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$. Formally, the relationship can be represented as:
(2) $\mathcal{E}(d)\approx\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]+\lambda w,\quad\lambda\in\mathbb{R}.$
Furthermore, when the stronger hypothesis assuming the clusters follow the Gaussian distribution is adopted, quantitative analysis results is derived in the appendix.
3.3. QAEncoder
Building upon the conical distribution hypothesis, we introduce a novel encoding strategy, $\text{QAE}_{\text{base}}$, which represents the document by the cluster center $\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$ of potential queries, instead of the document embedding $\mathcal{E}(d)$ itself. Formally, the transformation is defined as follows:
(3) $\text{QAE}_{\text{base}}(d)=\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]\approx\overline{\mathcal{E}(\mathcal{Q}(d))}=\frac{1}{n}\sum_{i=1}^{n}\mathcal{E}(q_{i}).$
@@ 3.3. QAEncoder
However, the ideal embedding model should not only bring related entries closer in the embedding space but also separate unrelated entries as much as possible. Despite the $\text{QAE}_{\text{base}}$ model enhances document-query relevance compared to the original document embedding $\mathcal{E}(d)$, it poses the distinguishability issue.
3.3.1. Document Fingerprint Strategy

### v3

@@ Abstract
Modern QA systems entail retrieval-augmented generation (RAG) for accurate and trustworthy responses. However, the inherent gap between user queries and relevant documents hinders precise matching. We introduce QAEncoder, a training-free approach to bridge this gap. Specifically, QAEncoder estimates the expectation of potential queries in the embedding space as a robust surrogate for the document embedding, and attaches document fingerprints to effectively distinguish these embeddings. Extensive experiments across diverse datasets, languages, and embedding models confirmed QAEncoder’s alignment capability, which offers a simple-yet-effective solution with zero additional index storage, retrieval latency, training costs, or catastrophic forgetting and hallucination issues. The repository is publicly available at https://github.com/IAAR-Shanghai/QAEncoder.
QAEncoder: Towards Aligned Representation Learning
in Question Answering Systems
Zhengren Wang1,2†, Qinhan Yu1†, Shida Wei1 Zhiyu Li2∗, Feiyu Xiong2, Xiaoxing Wang2, Simin Niu2, Hao Liang1,3, Wentao Zhang1,2,3∗ 1Peking University 2Institute for Advanced Algorithms Research 3Zhongguancun Academy wzr@stu.pku.edu.cn     {lizy,xiongfy}@iaar.ac.cn     wentao.zhang@pku.edu.cn
[…]
[1.5em]1.5em1em $\dagger$ Equal contribution; * Corresponding author.
1 Introduction
“What I cannot create, I do not understand."         — Richard Feynman
@@ 1 Introduction
Question Answering (QA) systems aim to generate accurate responses to user queries with applications in customer service (Xu et al., 2024), search engine (Ojokoh and Adebisi, 2018), healthcare (Guo et al., 2022) and education (Levonian et al., 2023). Modern QA systems leverage large language models (LLMs) such as ChatGPT (Achiam et al., 2023), supplemented with retrieval-augmented generation (RAG) to address issues of outdated or hallucinatory information, especially for rapidly evolving knowledge bases (Lewis et al., 2020; Huang et al., 2023; Gupta et al., 2024). The efficacy of RAG hinges on its retrieval module for identifying relevant documents from a vast corpus. Dense retrievers (Lewis et al., 2020; Hofstätter et al., 2021), contrasted with keyword-matching-based sparse retrievers (Jones, 1973; Robertson and Zaragoza, 2009), have enabled more precise retrieval by mapping queries and documents into a shared vector space. Despite advancements, a significant challenge that persists is bridging the gap between user queries and documents across lexical, syntactic, semantic, and content dimensions (Zheng et al., 2020; Nogueira et al., 2019; Cheriton, 2019), termed the document-query gap. Three main approaches have emerged to address this challenge: training-based, document-centric and query-centric alignment.
Training-based approaches (Dong et al., 2022; Li et al., 2022; W et al., 2023; Zhang et al., 2024a; Khanna and Subedi, 2024) directly train embedding models with QA datasets to close the representation of relevant queries and documents. But for out-of-domain or multi-domain adaption settings, they struggle to fully generalize (Suprem and Pu, 2022) and face catastrophic forgetting issues. Document-centric approaches (Wang et al., 2023b; Gao et al., 2023; Kim and Min, 2024) generate pseudo-documents for user queries via LLMs, which are then used as retrieval queries; however, this inference process is costly, time-consuming, and prone to hallucinations (Wang et al., 2023b).
In contrast, query-centric methods (Nogueira et al., 2019) generate and index potential queries to better match user queries. Yet, existing methods mainly focus on sparse retrievers and have not fully leveraged the potential of dense retrievers (Cheriton, 2019; Mallia et al., 2021). The naive attempt is to directly store predicted QA pairs into a vector database, but with evident drawbacks like expanded index size, longer retrieval latency, and limited query handling. Therefore, integrating query-centric methods with dense retrievers remains challenging and heavily under-explored.
Figure 1: Illustration of QAEncoder’s alignment process. Left: Solid lines represent diversified query generation, while dashed lines indicate Monte Carlo estimation. Right: The heatmap depicts the similarity scores among the embeddings of different queries, the document, and the mean estimation. Compared to the document itself, the mean estimation is significantly better aligned with different queries, i.e. a robust surrogate for the document embedding.
@@
Inspired by Feynman’s philosophy of learning, in QA systems, effective information retrieval extends beyond mere storage and involves creative processes like query formulation and summarization (Reyes et al., 2021). For instance, the well-established 5W1H framework (Who, What, When, Where, Why, How) help systematically deconstruct information and actively foster a deeper understanding (Jinks, 2019). In this paper, we continue the research line of query-centric methods, and propose QAEncoder as a pioneering work.
As demonstrated in Fig. 1, our method initially generates diversified queries (e.g. 5W1H), and then estimates the cluster center of potential queries by the Monte Carlo method. The similarity matrix reveals that, for any query, the mean-query similarity is significantly higher than both document-query and other query-query similarities. Hence, we advocate using the cluster center as a surrogate for the document embedding, which bridges the document-query gap robustly without extra index size and retrieval latency. Meanwhile, we propose the conical distribution hypothesis for more theoretical grounding, which intuitively visualizes the document-query gap and its alignment by abstracting geometric properties of the embedding space.
Despite these advantages, this basic proposal encounters a critical challenge. While enhancing similarity with user queries, it simultaneously reduces the distinguishability between document representations, as they all become query-like. To address this side effect, we further propose document fingerprint strategies to reintroduce unique document identities into representations and enable state-of-the-art performance.
@@
The contributions are threefold:
  • Methodological Innovations. We pioneer to bridge the document-query gap in dense retrievers from the query-centric perspective. Our method, QAEncoder, not only avoids extra index storage, retrieval latency, training cost and hallucination, but also guarantees diversified query handling and robust generalization. We also propose document fingerprints to address the side effect on distinguishability and achieve state-of-the-art performance.
  • Theoretical Discovery. We formulate the conical distribution hypothesis, and validate it through empirical analysis. This hypothesis provides not only deeper insights into the geometry of semantic space, but also theoretical foundations for the alignment process.
  • Practical Applications. QAEncoder is a simple-yet-effective plugin, which seamlessly integrates with existing RAG architectures and training-based methods. This integration significantly boosts system performance with minor modifications required.
2 Related Work
@@
Retrieval-augmented generation significantly improves large language models in QA systems by incorporating a retrieval module that fetches relevant information from external knowledge sources (Févry et al., 2020; Guu et al., 2020; Izacard and Grave, 2021; Zhao et al., 2024). Retrieval models have evolved from early sparse retrievers, such as TF-IDF (Jones, 1973) and BM25 (Robertson and Zaragoza, 2009), which rely on word statistics and inverted indices, to dense retrieval strategies (Lewis et al., 2020) that utilize neural representations for enhanced semantic matching. Advanced methods, such as Self-RAG (Asai et al., 2023) which determines if additional information is required and evaluates the relevance of retrieved content, and RAG-end2end (Siriwardhana et al., 2023) that jointly trains the retriever and generator, represent significant developments in this area. However, these methods still ignore the document-query gap.
@@
Training-based approaches bridge the document-query gap generally by contrastive learning (Xiong et al., ; Qu et al., 2021) or knowledge distillation (Zhang et al., 2024a; Khanna and Subedi, 2024). For instance, Dong et al. (2022) showed parameter sharing of the query encoder and the document encoder improves overall performance by projecting queries and documents into shared space. Dual-Cross-Encoder (Li et al., 2022) and Query-as-context (W et al., 2023) train embedding models from scratch with paired document-query samples. GPL (Wang et al., 2021), CAI (Iida and Okazaki, 2022) and AugTriever (Meng et al., 2022) conduct domain adaptation with pseudo-queries. However, training-based methods usually face generalization difficulty (Suprem and Pu, 2022), and catastrophic forgetting issues (Pan et al., 2024; Saunders, 2022).
@@
Document-centric methods, such as HyDE (Gao et al., 2023) and Query2doc (Wang et al., 2023b), dynamically transform user queries into pseudo-documents using LLMs for both sparse and dense retrievers. QA-RAG (Kim and Min, 2024) advances by implementing a two-way retrieval mechanism that utilizes both user query and pseudo-documents for respective retrieval and ranking. However, their effectiveness is highly dependent on the quality of pseudo-documents, which are prone to hallucinations, especially for the latest information. Furthermore, during inference, invoking LLMs for user queries imposes extra computational cost and latency, leading to degraded user experience.
@@
The seminal work, Doc2Query (Nogueira et al., 2019), focuses on the vocabulary mismatch problem for sparse retrievers by expanding the document with keywords in predicted queries. The improved version, DocT5Query (Cheriton, 2019), trains a T5 model to predict queries, but remains confined to sparse retrievers. There are also proposals to directly store question-answer pairs for retrieving and ranking, such as RePAQ (Lewis et al., 2021), QUADRo and QR (Campese et al., 2023, 2024). However, these attempts suffer from expanded index size, longer retrieval latency, and limited query handling. Hence, the query-centric alignment for dense retrievers remains challenging and largely unexplored.
@@
Figure 2: The conical distribution hypothesis. (a) t-SNE visualization of queries derived from various documents in the embedding space, illustrating distinct clustering behavior. (b) Angular distribution of document and query embeddings, showing the distribution of angles between $v_{d}=\mathcal{E}(d)-\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$ and $v_{q_{i}}=\mathcal{E}(q_{i})-\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$. The angles form a bell curve just below 90°, supporting that $v_{d}$ is approximately orthogonal to each $v_{q_{i}}$. (c) 3D visualization illustrating the conical distribution of the document (black point) and query (colored points) embeddings within a unit sphere. The star indicates the queries’ cluster center.
3 Method
3.1 Problem Formulation
Given a query $q$ and a document corpus $\mathcal{D}=\{d_{1},d_{2},...,d_{N}\}$, our task is to retrieve a subset of $K$ most relevant documents $\mathcal{D}_{+}=\{d_{i_{1}},d_{i_{2}},...,d_{i_{K}}\}$ through vector search. We define our embedding model as $\mathcal{E}(\cdot)$, which maps each document $d$ and query $q$ from the textual space $\mathcal{C}$ to a vector space $\mathbb{R}^{r}$. The semantic relevance is quantified by the cosine similarity, defined as: $sim(q,d)=\frac{\mathcal{E}(q)^{T}\mathcal{E}(d)}{\|\mathcal{E}(q)\|\|\mathcal{E}(d)\|}.$ Furthermore, for each document $d$ in our datasets, we invoke the query generator $\mathcal{Q}(\cdot)$ multiple times to generate $n$ predicted queries $\{q_{i}\}_{i=1}^{n}$, where the cluster center in embedding space is captured by $\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$ and Monte Carlo estimation $\overline{\mathcal{E}(\mathcal{Q}(d))}$.
3.2 QAEncoder
We first introduce a novel encoding method, $\text{QAE}_{\text{base}}$, which represents the document by the cluster center $\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$ of potential queries. Formally, the representation is defined as follows:
$\text{QAE}_{\text{base}}(d)$ $=\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$ (1)
$\approx\overline{\mathcal{E}(\mathcal{Q}(d))}=\frac{1}{n}\sum_{i=1}^{n}\mathcal{E}(q_{i}).$
@@ 3.2 QAEncoder
To thoroughly illustrate its mechanism, we formally define the conical distribution hypothesis and validate its reasonableness with empirical analysis.
@@
For any document $d$, the potential queries approximately form a single cluster on some hyperplane =$\mathcal{H}=\{x\in\mathbb{R}^{r}\mid w\cdot x=b\}$ in the semantic space, where $w\in\mathbb{R}^{r}$ is the normal vector and $b\in\mathbb{R}$ is the bias term. Furthermore, the document embedding $\mathcal{E}(d)$ lies on the perpendicular line intersecting the cluster center $\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]$. Formally, the relationship can be expressed as:
$\mathcal{E}(d)\approx\mathbb{E}[\mathcal{E}(\mathcal{Q}(d))]+\lambda w,\quad\lambda\in\mathbb{R}.$
We acknowledge that a more realistic model is an oblique cone, not the regular cone. Despite the hypothesis being highly simplified, it clearly captures the dynamics of QAEncoder’s alignment. As shown in Fig. 2, the document embedding $\mathcal{E}(d)$ is an outlier of the cluster of potential queries; whereas $\text{QAE}_{\text{base}}(d)$ bridges this substantial gap, mathematically denoted by $\lambda w$. We leave more discussions in Appendix A.2 for interested readers.
@@
Nonetheless, the ideal embedding model should not only bring related entries closer but also separate unrelated entries as much as possible. Despite $\text{QAE}_{\text{base}}$ increases document-query similarity, it poses the distinguishability issue.
3.2.1 Document Fingerprint Strategies
Figure 3: Architecture of QAEncoder. Left: Corpus documents are embedded using QAEncoder to obtain query-aligned representations for indexing. User queries are encoded with a vanilla encoder and used to retrieve relevant documents. Right: Internal mechanism of QAEncoder. QAEncoder addresses the document-query gap by generating a diverse set of queries for each document to create semantically aligned embeddings. Additionally, document fingerprint strategies are employed to ensure document distinguishability.